### v1

# Learning Content-enhanced Mask Transformer for Domain Generalized Urban-Scene Segmentation

Abstract
Domain-generalized urban-scene semantic segmentation (USSS) aims to learn generalized semantic predictions across diverse urban-scene styles. Unlike domain gap challenges, USSS is unique in that the semantic categories are often similar in different urban scenes, while the styles can vary significantly due to changes in urban landscapes, weather conditions, lighting, and other factors. Existing approaches typically rely on convolutional neural networks (CNNs) to learn the content of urban scenes.
In this paper, we propose a Content-enhanced Mask TransFormer (CMFormer) for domain-generalized USSS. The main idea is to enhance the focus of the fundamental component, the mask attention mechanism, in Transformer segmentation models on content information. To achieve this, we introduce a novel content-enhanced mask attention mechanism. It learns mask queries from both the image feature and its down-sampled counterpart, as lower-resolution image features usually contain more robust content information and are less sensitive to style variations. These features are fused into a Transformer decoder and integrated into a multi-resolution content-enhanced mask attention learning scheme.
Extensive experiments conducted on various domain-generalized urban-scene segmentation datasets demonstrate that the proposed CMFormer significantly outperforms existing CNN-based methods for domain-generalized semantic segmentation, achieving improvements of up to 14.00% in terms of mIoU (mean intersection over union). The source code for CMFormer will be made available at this repository.
1 Introduction
Urban-scene semantic segmentation (USSS) is a challenging problem because of the large scene variations due to changing landscape, weather, and lighting conditions [44, 31, 10, 4]. Unreliable USSS can pose a significant risk to road users. Hence, domain generalization is essential for robust USSS [35, 16, 7]. In contrast to common domain generalization, domain generalized USSS requires special attention because the domain gap is mainly caused by large style variations whereas changes in semantics largely remain consistent (see Fig. 1 for an example).
Existing approaches can be divided into two groups. One group focuses on the style de-coupling. This is usually achieved by a normalization [35, 16, 38] or whitening [36, 7, 57, 38] transformation. However, the de-coupling methodology falls short as the content is not learnt in a robust way. The other group is based on adverse domain training [65, 23, 66]. However, these methods usually do not particularly focus on urban styles and therefore their performance is limited.
In this paper, we focus on two main objectives for USSS: (1) a concise but expressive representation of the content, and (2) a robust module for style change handling. For the first goal, we propose to use Mask Transformer. Transformer segmentation models [47, 5, 51, 60, 61, 6] provide more expressive representations than CNN-based methods because of their mask attention mechanisms allowing to incorporate long-range dependencies [27, 28, 11], context [19], and saliency [5, 6]. For the second goal, mask-level learning is used which is less sensitive than pixel-level learning in handling style changes [34, 12, 52].
Figure 1: (a) Example of domain generalized USSS (top: source domain, rest: target domain); (b) The key challenge lies in the domain gap is mainly from the extremely-varied styles but the content is similar; (c) The proposed CMFormer with content-enhanced mask attention focus on: 1) learn an expressive content representation, 2) robust to style variation. SAW: the present state-of-the-art [38].
To this end, a novel Content-enhanced Mask TransFormer (CMFormer) is proposed. Particularly, a content-enhanced mask attention mechanism is introduced, which takes the original image feature together with its down-sampled counterpart as input. The down-sampled features contain more content and are less sensitive to style variations. Both features are fused to learn the context. Then, the content-enhanced mask attention is extended to incorporate multi-resolution features.
Large-scale experiments are conducted with various domain generalized USSS settings, i.e., trained on one dataset from [42, 43, 8, 33, 59] as the source domain, and validated on the rest of the four datasets as the unseen target domains. The experiments show that the proposed CMFormer achieves upto 14.00% mIoU improvement compared to the state-of-the-art (e.g., SAW [38], WildNet [23]). The proposed CMFormer also outperforms existing mask-level Transformer segmentation methods and the generic segmentation methods pre-trained by the foundation model (i.e., SAM [21], SegGPT [54]). It also shows state-of-the-art performance on synthetic-to-real and ideal-to-adverse generalization.
Our contribution is summarized as follows:
  • To the best of our knowledge, it is the first attempt to use ViT for the task of domain generalized USSS.
  • A content-enhanced mask attention mechanism is proposed enhancing the content information from image features.
  • A multi-resolution content-enhanced mask attention learning scheme is proposed for expressive content representation and style robustness.
  • Extensive experiments show a large performance improvement over existing SOTA by upto 14.00% mIoU.
2 Related Work
Domain Generalization has been studied on no task-specific scenarios in the field of both machine learning and computer vision. Harary et al. [13] considered domain generalization in an unsupervised manner by learning a domain bridge. Hu et al. [15] proposed a framework for image retrieval in an unsupervised setting. Zhou et al. [67] proposed a framework to generalize to new homogeneous domains. Qiao et al. [41] and Peng et al. [39] proposed to learn domain generalization from a single source domain. Many other methods have been proposed such as entropy regularization [64], casual matching [29], extrinsic-intrinsic interaction [53], balance invariance [2], batch normalization embeddings [45] and multiple latent domain modeling [30].
Domain Generalized Semantic Segmentation can be regarded as the exploration of the prior unsupervised domain adaption segmentation task [38, 37, 63], but requires a larger generalization capability of a model on a variety of target domains. Existing methods focus on the generalization of in-the-wild [40], scribble [48] and multi-source images [20, 22], where substantial alterations can occur in both the content and style.
Domain Generalized USSS focuses on the generalization of driving-scenes [8, 59, 33, 43, 42]. These methods use either a normalization transformation (e.g., IBN [35], IN [16], SAN [38]) or whitening transformation (e.g., IW [36], ISW [7], DIRL [57], SAW [38]) on the training domain, to enable the model to generalize better on the target domains. Other advanced methods for domain generalization in segmentation typically rely on external images to incorporate more diverse styles [23, 65, 66, 25], and leverage content consistency across multi-scale features [62]. To the best of our knowledge, all of these methods are based on CNN.
Mask Transformer for Semantic Segmentation Earlier ViT based segmentation models (e.g., SegFormer [56]) follow the CNN based pipeline. Recently, the focus is on mask-level pipelines, which use the queries in the Transformer decoder to learn the masks, e.g., Segmenter [47], MaskFormer [5], Max-DeepLab [51], CMT-DeepLab [60], kMaX-DeepLab [61] and etc. More recently, Mask2Former [6] further simplified the pipeline of MaskFormer and achieves better performance.
3 Methodology
3.1 Problem Definition
Domain Generalization can be formulated as a worst-case problem [24, 66, 50]. Given a source domain $\mathcal{S}$, and a set of unseen target domains $\mathcal{T}_{1},\mathcal{T}_{2},\cdots$, a model parameterized by $\theta$ with the task-specific loss $\mathcal{L}_{task}$, the generic domain generalization task can be formulated as a worst-case problem, given by
$\mathop{\rm{min}}\limits_{\theta}\mathop{\rm{supp}}\limits_{\mathcal{T}:D(\mathcal{S};\mathcal{T}_{1},\mathcal{T}_{2},\cdots)\leq\rho}\mathbb{E}_{T}[\mathcal{L}_{task}(\theta;\mathcal{T}_{1},\mathcal{T}_{2},\cdots)],$ (1)
where $\theta$ denotes the model parameters, $D(\mathcal{S};\mathcal{T}_{1},\mathcal{T}_{2},\cdots)$ corresponds to the distance between the source $\mathcal{S}$ and target domain $\mathcal{T}$, and $\rho$ denotes the constraint threshold.
Domain generalized USSS is challenging as it has similar content (e.g., semantics) but variation of
Figure 2: (a) Feature domain for urban-scenes, where the content is similar while the styles have clear gaps; (b) The goal is robust style handing and stable content representing.
style (e.g., urban landscapes, weather, season, illumination) among each domain $\mathcal{S}$, $\mathcal{T}_{1}$, $\mathcal{T}_{2}$, $\cdots$. Visual examples have been provided in Fig. 1. Here we analyze the feature domain. As shown in Fig. 2a, while the content distribution is similar, the style distribution is separated and accounts for the domain gap. Our goal is to find a representation where both style and content are similarly distributed among domains, therefore minimizing the domain gap (in Fig.2b).
3.2 Content-enhanced Mask Attention
Figure 3: (a) The proposed Content-enhanced Mask Attention (CMA). Apart from the conventional mask attention (left branch), the down-sampled feature maps, which contain more abundant content information, are used to learn content enhanced query features (right branch). (b) Framework overview of the proposed Content-enhanced Mask TransFormer (CMFormer) for domain generalized semantic segmentation. The image decoder is directly inherited from the Mask2Former [6].
Our proposed idea is to enhance the focus of the mask attention mechanism [5, 6] on content information from urban scenes. This enhancement aids the segmentation masks in concentrating on scene content while reducing sensitivity to style variations. Since content/style information is usually more pronounced in the low/high resolution features of an image, it is intuitive to amplify the influence of low-resolution features in the self-attention learning.
Mask Attention Mechanism learns the query features as the segmentation masks by introducing a mask attention matrix based on the self-attention mechanism. Let $\mathbf{F}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ denote the image features from the image decoder with a spatial width and height of $W_{l}$ and $H_{l}$, and let $\mathbf{X}_{l}\in\mathbb{R}^{N\times C}$ denote the features of the $l^{th}$ layer in a Transformer decoder, where $N$ is the number of semantic categories. Further, $C_{F}$ and $C$ denote the channel dimensions of the image features and self-attention embeddings respectively. $\mathbf{X}_{0}$ refers to the input query features of the Transformer decoder. The key $\mathbf{K}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C}$, value $\mathbf{V}_{l}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C}$, and query $\mathbf{Q}_{l}\in\mathbb{R}^{N\times C}$ is computed by linear transformations $f_{K}$, $f_{V}$ and $f_{Q}$, respectively, given by
$\mathbf{K}_{l}=f_{K}(\mathbf{F}_{l-1}),\mathbf{V}_{l}=f_{V}(\mathbf{F}_{l-1}),\mathbf{Q}_{l}=f_{Q}(\mathbf{X}_{l-1}).$ (2)
Then, the query feature $\mathbf{X}_{l}$ is computed by
$\mathbf{X}_{l}={\rm softmax}(\mathcal{M}_{l-1}+\mathbf{Q}_{l}\mathbf{K}_{l}^{\mathsf{T}})\mathbf{V}_{l}+\mathbf{X}_{l-1},$ (3)
where $\mathcal{M}_{l-1}\in\{0,1\}^{N\times H_{l}W_{l}}$ is a binary mask attention matrix from the resized mask prediction of the previous $(l-1)^{th}$ layer, with a threshold of 0.5. $\mathcal{M}_{0}$ is binarized and resized from $\mathbf{X}_{0}$. It filters the foreground regions of an image, given by
$\mathcal{M}_{l-1}(x,y)=\begin{cases}0&\text{if $\mathcal{M}_{l-1}(x,y)=1$ }\\ $ - $\infty$ $&\text{else}\end{cases}.$ (4)
Content Enhancement is incorporated into the mask attention mechanism by acquiring supplementary query features from low-resolution image features. This is based on the understanding that content information tends to be more discriminative, while style information is less sensitive in low-resolution image features. To this end, a parallel branch that leverages low-resolution image features is introduced in each Transformer decoder layer (in Fig. 3a). The content-enhanced image feature $\mathbf{F}_{l}^{c}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ is computed by average pooling $\rm{avgpool}$ from the original image feature $\mathbf{F}_{l}^{c}\in\mathbb{R}^{(W_{l}\cdot H_{l})\times C_{F}}$ by
$\mathbf{F}_{l}^{c}={\rm avgpool}(\mathbf{F}_{l}).$ (5)
Following Eq. 2, for the $l^{th}$ Transformer decoder, the content enhanced key $\mathbf{K}^{c}_{l}\in\mathbb{R}^{(W_{l}/2\cdot H_{l}/2)\times C}$ and value $\mathbf{V}^{c}_{l}\in\mathbb{R}^{(W_{l}/2\cdot H_{l}/2)\times C}$ are computed from the linear transformation of the down-sampled image feature $\mathbf{F}_{l}^{c}$. Then, similar to Eq. 3, the content-enhanced query feature $\mathbf{X}_{l}^{c}$ is computed as
$\mathbf{X}_{l}^{c}={\rm softmax}(\mathcal{M}_{l-1}^{c}+\mathbf{Q}_{l}\mathbf{K}_{l}^{c\mathsf{T}})\mathbf{V}_{l}^{c}+\mathbf{X}_{l-1}^{c},$ (6)
where $\mathcal{M}_{l-1}^{c}\in\{0,1\}^{N\times(H_{l}/2\cdot W_{l}/2)}$ also follows the computation of $\mathcal{M}_{l-1}$ defined in Eq. 4.
Content-aware Query Feature Fusion is proposed to merge the query $\mathbf{X}_{l}$ with content-enhanced query features $\mathbf{X}_{l}^{c}$, allowing for the highlighting of content while retaining the necessary details for semantic prediction. The fused feature $\mathbf{X}_{l}^{final}$ serves as the final output of the $l^{th}$ Transformer decoder, and it is computed as follows:
$\mathbf{X}_{l}^{final}=h_{l}([\mathbf{X}_{l},\mathbf{X}_{l}^{c}]),$ (7)
Here, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] represents the concatenation operation, and $h_{l}(\cdot)$ refers to a linear layer.
3.3 Content-enhancement for Multi-resolution Features
In the process of decoding domain-generalized segmentation predictions from coarse-to-fine image features, it is crucial to enhance the content information at each resolution. This ensures that the content remains highlighted in the query features of masks, allowing for the learning of segmentation masks that are less sensitive to style variations.
Multi-resolution Features include $\times$32, $\times$16 and $\times$8 resolution features from the image decoder, denoted as $\mathbf{F}^{\times 32}$, $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$, respectively. Following Eq. 23, the key, value and query for $\mathbf{F}^{\times 32}$ and $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$ are denoted by $\mathbf{K}^{\times 32}$, $\mathbf{V}^{\times 32}$, $\mathbf{Q}^{\times 32}$ and $\mathbf{K}^{\times 16}$, $\mathbf{V}^{\times 16}$, $\mathbf{Q}^{\times 16}$ and $\mathbf{K}^{\times 8}$, $\mathbf{V}^{\times 8}$, $\mathbf{Q}^{\times 8}$, respectively.
Content-enhanced Multi-resolution Features are down-sampled from $\times$32, $\times$16 and $\times$8 resolution features $\mathbf{F}^{\times 32}$ and $\mathbf{F}^{\times 16}$ and $\mathbf{F}^{\times 8}$ into $\mathbf{F}^{\times 32,c}$ and $\mathbf{F}^{\times 16,c}$ and $\mathbf{F}^{\times 8,c}$ by following Eq. 5. Also following Eq. 23, the key, value and query for $\mathbf{F}^{\times 32,c}$ and $\mathbf{F}^{\times 16,c}$ and $\mathbf{F}^{\times 8,c}$ are denoted as $\mathbf{K}^{\times 32,c}$, $\mathbf{V}^{\times 32,c}$, $\mathbf{Q}^{\times 32,c}$ and $\mathbf{K}^{\times 16,c}$, $\mathbf{V}^{\times 16,c}$, $\mathbf{Q}^{\times 16,c}$ and $\mathbf{K}^{\times 8,c}$, $\mathbf{V}^{\times 8,c}$, $\mathbf{Q}^{\times 8,c}$, respectively.
Multi-resolution Feature Fusion to the Transformer Decoder directly follows the original Mask2Former [6], which alternatively learns the image features from each different resolution. As there are 9 self-attention layers in the Transformer decoder, the first, forth, seventh layers take the ×$\times 32$ image features as input (black down-arrows in Fig. 3b), the second, fifth, eighth layers take the ×$\times 16$ image features as the input (blue downarrows in Fig. 3b), and the third, sixth, ninth layer take the ×$\times 8$ image features as the input (green downarrows in Fig. 3b).
Taking the first, forth, seventh layers with ×$\times 32$ image features as an example (here $l=1,4,7$), the content-enhanced learning process is given by
$\mathbf{X}_{l}={\rm softmax}(\mathcal{M}_{l-1}+\mathbf{Q}_{l}^{\times 32}\mathbf{K}_{l}^{\times 32\mathsf{T}})\mathbf{V}_{l}^{\times 32}+\mathbf{X}_{l-1}^{final},$ (8)
$\mathbf{X}_{l}^{c}={\rm softmax}(\mathcal{M}_{l-1}^{c}+\mathbf{Q}_{l-1}^{\times 32,c}\mathbf{K}^{\times 32,c\mathsf{T}})\mathbf{V}^{\times 32,c}+\mathbf{X}_{l-1}^{final},$ (9)
Following Eq. 7, $\mathbf{X}_{l}^{c}$ and $\mathbf{X}_{l}$ are merged to compute the final output of this decoder layer, denoted as $\mathbf{X}_{l}^{final}$. For the first layer, $\mathbf{X}_{0}^{final}$ is equivalent to $\mathbf{X}_{0}$, which is the input query features of the Transformer decoder.
The content enhance learning of the second, fifth and eighth layer follows Eq. 89 but uses the $\mathbf{K}^{\times 16,c}$, $\mathbf{V}^{\times 16,c}$, $\mathbf{Q}^{\times 16,c}$ and $\mathbf{K}^{\times 16}$, $\mathbf{V}^{\times 16}$, $\mathbf{Q}^{\times 16}$ as input. The third, sixth and ninth layer also follow Eq. 89 but takes $\mathbf{K}^{\times 8,c}$, $\mathbf{V}^{\times 8,c}$, $\mathbf{Q}^{\times 8,c}$ and $\mathbf{K}^{\times 8}$, $\mathbf{V}^{\times 8}$, $\mathbf{Q}^{\times 8}$ as input.
3.4 Network Architecture and Implementation Details
The overall framework is shown in Fig. 3b. The Swin-Base Transformer [27] is used as the backbone, and the image decoder is inherited from Mask2Former [6]. The pre-trained model from ImageNet [9] is utilized for initialization.
The image decoder from [6] uses the off-the-shelf multi-scale deformable attention Transformer (MSDeformAttn) [68] with the default setting in [68, 6]. By considering the image features from the Swin-Based encoder as input, every 6 MSDeformAttn layers are used to progressively up-sample the image features in ×$\times 32$, ×$\times 16$, ×$\times 8$, and ×$\times 4$, respectively. The $1/4$ resolution feature map is fused with the features from the Transformer decoder for dense prediction.
Following the default setting of MaskFormer [5] and Mask2Former [6], the final loss function $\mathcal{L}$ is a linear combination of the binary cross-entropy loss $\mathcal{L}_{ce}$, dice loss $\mathcal{L}_{dice}$, and the classification loss $\mathcal{L}_{cls}$, given by
$\mathcal{L}=\lambda_{ce}\mathcal{L}_{ce}+\lambda_{dice}\mathcal{L}_{dice}+\lambda_{cls}\mathcal{L}_{cls},$ (10)
with hyper-parameters $\lambda_{ce}=\lambda_{dice}=5.0,\lambda_{cls}=2.0$ as the default of Mask2Former without any tuning. The Adam optimizer is used with an initial learning rate of $1\times 10^{-4}$. The weight decay is set 0.05. The training terminates after 50 epochs.
4 Experiment
4.1 Dataset & Evaluation Protocols
Building upon prior research in domain-generalized USSS, our experiments utilize five different semantic segmentation datasets. Specifically, CityScapes [8] provides 2,975 and 500 well-annotated samples for training and validation, respectively. These driving-scenes are captured in Germany cities with a resolution of 2048$\times$1024. BDD-100K [59] also provides diverse urban driving scenes with a resolution of 1280$\times$720. 7,000 and 1,000 fine-annotated samples are provided for training and validation of semantic segmentation, respectively. Mapillary [33] is also a real-scene large-scale semantic segmentation dataset with 25,000 samples. SYNTHIA [43] is large-scale synthetic dataset, and provides 9,400 images with a resolution of 1280$\times$760. GTA5 [42] is a synthetic semantic segmentation dataset rendered by the GTAV game engine. It provides 24,966 simulated urban-street samples with a resolution of 1914$\times$1052. We use C, B, M, S and G to denote these five datasets.
Following prior domain generalized USSS works [35, 36, 7, 38], the segmentation model is trained on one dataset as the source domain, and is validated on the rest of the four datasets as the target domains. Three settings include: 1) G to C, B, M, S; 2) S to C, B, M, G; and 3) C to B, M, G, S. mIoU (in percentage %) is used as the validation metric. All of our experiments are performed three times and averaged for fair comparison. All the reported performance is directly cited from prior works under the ResNet-50 backbone [35, 36, 7, 38].
4.2 Comparison with State-of-the-art Domain Generalized USSS Methods
[…]
Method Proc. & Year Trained on GTA5 (G) Trained on SYNTHIA (S) Trained on Cityscapes (C)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ S $\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
IBN [35] ECCV2018 33.85 32.30 37.75 27.90 32.04 30.57 32.16 26.90 48.56 57.04 45.06 26.14
IW [36] CVPR2019 29.91 27.48 29.71 27.61 28.16 27.12 26.31 26.51 48.49 55.82 44.87 26.10
Iternorm [17] CVPR2019 31.81 32.70 33.88 27.07 - - - - 49.23 56.26 45.73 25.98
DRPC [62] ICCV2019 37.42 32.14 34.12 28.06 35.65 31.53 32.74 28.75 49.86 56.34 45.62 26.58
ISW [7] CVPR2021 36.58 35.20 40.33 28.30 35.83 31.62 30.84 27.68 50.73 58.64 45.00 26.20
GTR [37] TIP2021 37.53 33.75 34.52 28.17 36.84 32.02 32.89 28.02 50.75 57.16 45.79 26.47
DIRL [57] AAAI2022 41.04 39.15 41.60 - - - - - 51.80 - 46.52 26.50
SHADE [65] ECCV2022 44.65 39.28 43.34 - - - - - 50.95 60.67 48.61 27.62
SAW [38] CVPR2022 39.75 37.34 41.86 30.79 38.92 35.24 34.52 29.16 52.95 59.81 47.28 28.32
WildNet [23] CVPR2022 44.62 38.42 46.09 31.34 - - - - 50.94 58.79 47.01 27.95
AdvStyle [66] NeurIPS2022 39.62 35.54 37.00 - 37.59 27.45 31.76 - - - - -
CMFormer (Ours) 2023 55.31 49.91 60.09 43.80 44.59 33.44 43.25 40.65 59.27 71.10 58.11 40.43
$\uparrow$10.66 $\uparrow$10.63 $\uparrow$14.00 $\uparrow$12.46 $\uparrow$5.67 $\downarrow$1.80 $\uparrow$8.73 $\uparrow$11.49 $\uparrow$6.32 $\uparrow$10.43 $\uparrow$ 9.50 $\uparrow$12.11
GTA5 Source Domain The third column of Table 1 reports the performance on target domains of C, B, M and S, respectively. The proposed CMFormer shows a performance improvement of 10.66%, 10.63%, 14.00% and 12.46% compared to existing state-of-the-art CNN based methods on each target domain, respectively. These outcomes demonstrate the feature generalization ability of the proposed CMFormer. Notice that the source domain GTA5 is a synthetic dataset, while the target domains are real images. It further validates the performance of the proposed method.
SYNTHIA Source Domain The forth column of Table 1 reports the performance. The proposed CMFormer shows a 5.67%, 8.73% and 11.49% mIoU performance gain against the best CNN based methods, respectively. However, on the BBD-100K (B) dataset, the semantic-aware whitening (SAW) method [38] outperforms the proposed CMFormer by 1.80% mIoU. Nevertheless, the proposed CMFormer still outperforms the rest state-of-the-art methods. The performance gain of the proposed CMFormer when trained on SYNTHIA dataset is not as significant as it is trained on CityScapes or GTA5 dataset. The explanation may be that the SYNTHIA dataset has much fewer samples than GTA5 dataset, i.e., 9400 v.s. 24966, and a transformer may be under-trained.
CityScapes Source Domain The last column of Table 1 reports the performance. The proposed CMFormer shows a performance gain of 6.32%, 10.43%, 9.50% and 12.11% mIoU on the B, M, G and S dataset against the state-of-the-art CNN based method. As BDD100K dataset contains many nigh-time urban-street images, it is particularly challenging for existing domain generalized USSS methods. Still, a performance gain of 6.32% is observed by the proposed CMFormer.
4.3 Ablation Studies
[…]
Table 2: Ablation studies on each component of the proposed CMFormer. $\times$32, $\times$16 and $\times$8 denote the image features of $\times$32, $\times$16 and $\times$8 resolution. $\checkmark$ refers to the content enhancement is implemented. Evaluation metric mIoU is presented in (%).
Table 3: Ablation studies on the embedding order of the multi-resolution features in the CMFormer. Ours are compared with the plain and descent settings. Evaluation metric mIoU in %.
Content Enhancement Trained on CityScapes (C) Trained on SYNTHIA (S)
$\times$32 $\times$16 $\times$8 $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S $\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G
55.43 66.12 55.05 38.19 42.64 31.91 36.80 37.29
$\checkmark$ 56.17 67.55 55.42 38.83 42.77 31.62 41.61 37.56
$\checkmark$ $\checkmark$ 58.10 69.72 55.54 39.41 44.22 33.19 42.23 38.78
$\checkmark$ $\checkmark$ $\checkmark$ 59.27 71.10 58.11 40.43 44.59 33.44 43.25 40.65
Method Trained on Cityscapes (C)
$\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
$plain$ 59.20 71.08 57.89 40.16
$descent$ 58.51 71.01 57.53 40.29
Ours 59.27 71.10 58.11 40.43
Table 3: Ablation studies on the embedding order of the multi-resolution features in the CMFormer. Ours are compared with the plain and descent settings. Evaluation metric mIoU in %.
To better understand the content enhancement for mask attention, some segmentation masks learnt by the existing mask attention and our content enhanced mask attention are visualized in the second and third row of Fig. 4. The proposed content enhanced mask attention is less sensitive to the style variation. The learnt masks are more capable to separate the key objects from the background.
Figure 4: Visualization of the segmentation masks learnt by the original mask attention (in second row) and the proposed content-enhanced mask attention (in third row), under the C [8] $\rightarrow$ B [59], G [42], M [33], S [43] settings (in first row). The proposed content-enhanced mask attention is less sensitive to the style variation and is of better capability to separate the objects from background.
On Multi-resolution Feature Embedding Strategy is inherited from Mask2Former in an alternative manner [6]. To investigate its effectiveness, the method is compared with two alternative solutions: (1) plain sequence embedding (denoted as $plain$) and (2) alternative embedding in a descent order (denoted as $descent$). For the plain sequence embedding, the first, second and third three decoder layers are fed into the ×$\times 32$, ×$\times 16$ and ×$\times 8$ image features, respectively. For the alternative embedding in a descent order, each three decoder layers are fed into the image features with the order of ×$\times 8$, ×$\times 16$ and ×$\times 32$. The results are listed in Table 3. The embedding strategy inherited from [6] has a slightly better performance compared to the two other solutions.
4.4 Comparison with Recent State-of-the-art in Related Tasks
The proposed CMFormer is compared against three categories of recent state-of-the-art approaches that are relevant to related tasks. While these approaches do not specifically focus on domain-generalized USSS, we compare them to further highlight the optimality of the proposed CMFormer.
Specially, we consider mask-level Transformer segmentation models (Segmenter [47], MaskFormer [5], Mask2Former [6], denoted as $\mathcal{M}$), combining state-of-the-art CNN based domain generalization techniques with segmentation Transformer (denoted as $\mathcal{G}$), and the plain segmentation Transformer that is pre-trained from the foundation models (SAM [21], SegGPT [54], denoted as $\mathcal{F}$). It can be seen that the proposed CMFormer significantly outperforms other methods. It is also very interesting to observe that the whitening transformation (IW [36], ISW [36], SAW [38]) for CNN based domain generalized segmentation frameworks do not work for mask-level Transformer segmentation pipeline.
[…]
Method Backbone Category Trained on Cityscapes (C)
$\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
Segmenter [47] ViT-B $\mathcal{M}$ 31.42 37.76 28.56 17.32
Max-DeepLab [51] MaX-L 47.82 54.05 39.82 27.15
CMT-DeepLab [60] Axial-R104 48.74 57.93 41.66 26.06
kMaX-DeepLab [61] ConvNeXt-B 50.58 59.06 43.92 28.70
MaskFormer [5] Swin-B 51.86 64.70 49.24 33.35
Mask2Former [6] Swin-B 55.43 66.12 55.05 38.19
IN [35] + Mask2Former [6] Swin-B $\mathcal{G}$ 56.78 67.97 56.07 40.16
IW [36] + Mask2Former [6] Swin-B 54.20 63.14 50.57 35.59
ISW [36] + Mask2Former [6] Swin-B 53.00 63.16 52.14 36.97
SAW [38] + Mask2Former [6] Swin-B 55.78 62.36 57.88 39.96
SAM Pre-trained [21] ViT-H $\mathcal{F}$ 52.36 65.72 51.49 34.25
SegGPT Pre-trained [54] ViT-L 50.94 61.37 44.85 32.36
CMFormer (Ours) Swin-B $\mathcal{G}$ 59.27 71.10 58.11 40.43
4.5 Extra Experiments on Generalization
To Adverse Domains Beyond the standard validation protocols in existing methods [35, 16, 36, 7, 38, 66], we further validate the proposed CMFormer’s performance by benchmarking it on the adverse conditions dataset with correspondance (ACDC) [44]. It is the largest semantic segmentation dataset under a variety of adverse conditions, including rain, fog, night and snow. We set the fog, night, rain and snow as four different unseen domains, and directly use the model pre-trained on CityScapes for inference. The results are shown in Table 6. It significantly outperforms existing domain generalized segmentation methods (IN [35], Iternorm [16], IW [36], ISW [7], ISSA [25]) by upto 10.3%, 0.5%, 11.6%, 11.1% on the fog, night, rain and snow domains, respectively. On the other hand, it shows a superior performance of 4.4%, 4.0% and 1.8% against the second best-performed method Mask2Former on the fog, rain and snow domain, respectively.
[…]
Method Backbone Trained on Cityscapes (C)
$\rightarrow$ Fog $\rightarrow$ Night $\rightarrow$ Rain $\rightarrow$ Snow mean
IN [35] ResNet-50 63.8 21.2 50.4 49.6 43.7
Iternorm [16] ResNet-50 63.3 23.8 50.1 49.9 45.3
IW [36] ResNet-50 62.4 21.8 52.4 47.6 46.6
ISW [7] ResNet-50 64.3 24.3 56.0 49.8 48.1
ISSA [25] MiT-B2 67.5 33.2 55.9 53.2 52.5
RefineNet [26] ResNet-101 46.4 29.0 52.6 43.3 43.7
DeepLabv3+ [3] ResNet-101 45.7 25.0 50.0 42.0 41.6
DANet [12] ResNet-101 34.7 19.1 41.5 33.3 33.1
HRNet [52] HR-w48 38.4 20.6 44.8 35.1 35.3
SegFormer [56] MiT-B2 59.2 38.9 62.5 58.2 56.2
Mask2Former [6] Swin-B 73.4 37.1 63.6 62.5 58.0
CMFormer (Ours) Swin-B 77.8 33.7 67.6 64.3 60.1
Backbone Trained on Synthetic Domain (G+S)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M mean
ResNet-50 35.46 25.09 31.94 30.83
IBN [35] 35.55 32.18 38.09 35.27
ISW [7] 37.69 34.09 38.49 36.75
SHADE [65] 47.43 40.30 47.60 45.11
AdvStyle [66] 39.29 39.26 41.14 39.90
Ours 59.70 53.36 61.61 58.22
[…]
4.6 Visualization
T-SNE visualization To better understand the style change handling and stable robust representation learning, the feature space with samples when using B, M, G, S as target domains are shown in Fig. 5 by t-SNE. ISW [7], a typical style-decoupling method, has some difficulty to handle the style variation, but the distance of the semantic embeddings from each domain is close. Advstyle [66], a typical adverse style augmentation method, can better mix the styles, but the distance of semantic embeddings is not properly reduced. The original mask attention from [6] can reduce the distance of each semantic embedding but still has some difficulty to mix the different styles. In contrast, the proposed CMFormer can not only reduce the distance of semantic embeddings, but also allow the styles to be more uniformly distributed and therefore minimizes the domain gap.
Figure 5: T-SNE visualization of the features space on target domains. (a) ISW [7], (b) AdvStyle [66], (c) Mask2Former [6], (d) the proposed CMFormer. Our method has the most uniform distribution, which means the domain gap is minimized.
Quantitative segmentation prediction Some generalized segmentation prediction results on the C $\rightarrow$ B, M, G, S setting (in Fig. 9) and on the C $\rightarrow$ adverse domain setting (in Fig. 10). Compared with the CNN based methods, the proposed CMFormer shows a better segmentation prediction, especially in terms of the completeness of objects.
[…]
Figure 7: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN [35], IW [36], Iternorm [17], ISW [7]) and the proposed CMFormer under the C $\rightarrow$ adverse domain setting. More prediction is in the supplementary material.
5 Conclusion
In this paper, we explored the feasibility of adapting the mask Transformer for domain-generalized urban-scene semantic segmentation (USSS). To address the challenges of style variation and robust content representation, we proposed a content-enhanced mask attention mechanism. This mechanism is designed to capture more resilient content features while being less sensitive to style variations. Furthermore, we extended it to incorporate multi-resolution features and integrate it into a novel framework called the Content-enhanced Mask TransFormer (CMFormer). To evaluate the effectiveness of CMFormer, we conducted extensive experiments on multiple settings. The results demonstrated the superior performance of CMFormer compared to existing domain-generalized USSS methods.
Limitation Discussion & Boarder Social Impact. The proposed content-enhanced mask attention mechanism is derived from the self-attention mechanism and can be seamlessly integrated into ViT based segmentation pipelines. The experimental results demonstrate its superior performance, thereby indicating the potential to shift the focus of domain-generalized USSS towards ViT based pipelines.
Given the criticality of safety-related road applications, the proposed method holds significant importance. It has the potential to enhance the accuracy and reliability of semantic segmentation models, thereby contributing to safer and more efficient autonomous systems. Overall, the proposed content-enhanced mask attention mechanism not only offers promising advancements in domain-generalized USSS but also holds potential for broader applications in real-world scenarios.